# MASQUE: A Text-Guided Diffusion-Based Framework for Localized and Customized Adversarial Makeup

### v1

@@ Abstract
As facial recognition is increasingly adopted for government and commercial services, its potential misuse has raised serious concerns about privacy and civil rights. To counteract, various anti-facial recognition techniques have been proposed for privacy protection by adversarially perturbing face images, among which generative makeup-based approaches are the most popular. However, these methods, designed primarily to impersonate specific target identities, can only achieve weak dodging success rates while increasing the risk of targeted abuse. In addition, they often introduce global visual artifacts or a lack of adaptability to accommodate diverse makeup prompts, compromising user satisfaction. To address the above limitations, we develop MASQUE, a novel diffusion-based framework that generates localized adversarial makeups guided by user-defined text prompts. Built upon precise null-text inversion, customized cross-attention fusion with masking, and a pairwise adversarial guidance mechanism using images of the same individual, MASQUE achieves robust dodging performance without requiring any external identity. Comprehensive evaluations on open-source facial recognition models and commercial APIs demonstrate that MASQUE significantly improves dodging success rates over all baselines, along with higher perceptual fidelity and stronger adaptability to various text makeup prompts.
@@ 1 Introduction
Facial recognition (FR) systems [22] have been adopted in a wide range of security, biometrics, and commercial applications. However, their unregulated deployment poses serious privacy risks, allowing for unauthorized surveillance and malicious tracking. To address these concerns, anti-facial recognition (AFR) technologies [34] have emerged to protect user privacy from unauthorized FR systems. AFR techniques vary depending on which stage of facial recognition they disrupt and typically work by modifying images before they are shared online. Among them, adversarial methods are particularly effective, subtly altering face images to evade detection while preserving their natural appearance. Traditional adversarial approaches, such as noise-based methods [13, 37, 41], obscure facial features with norm-bounded global perturbations, while patch-based techniques [36, 15] optimize adversarial patterns in localized image regions. However, these methods often introduce noticeable visual artifacts, compromising the usability of the produced images.
@@
To overcome these limitations, recent work has shifted toward generative approaches, leveraging generative adversarial nets or diffusion models for AFR. Instead of relying on pixel-wise perturbations, these methods generate unrestricted yet semantically meaningful modifications that blend naturally into facial features. A notable direction is makeup-based AFR [38, 12, 26, 30, 7], which seamlessly integrates adversarial perturbations into makeup—a plausible approach, as makeup is inherently associated with facial appearance (see Appendix A for detailed discussions of related work).
@@
While makeup-based AFR methods offer a promising balance between privacy protection and aesthetics, they often struggle to preserve fine-grained facial details or fully adhere to user instructions from diverse prompts. In addition, these methods require images of an external target identity to guide the generation process for adversarial makeup transfer and primarily focus on the impersonation setting for evaluation, increasing the risks of targeted abuse. When considering the more privacy-critical dodging scenarios, their performance in protection success rates also drops significantly.
@@
Contribution. We propose MASQUE, a novel diffusion-based image editing method for localized adversarial makeup generation with customized text guidance. In particular, MASQUE is designed to meet the following desiderata simultaneously:
  • Inoffensive Identity Protection: ensure high protection success rates in dodging scenarios without requiring images of a specific target identity.
  • Localized Modification: restrict adversarial perturbations to designated areas, preserving all other facial details.
  • User Control: demonstrate strong prompt-following ability and adaptability to diverse text makeup prompts.
@@ 2 Problem Formulation
In this section, we specify the threat model for unauthorized facial recognition (Section 2.1), introduce the background on anti-facial recognition (Section 2.2), and explain why dodging is more suitable than impersonation (Section 2.3).
2.1 Threat Model
We consider the adversarial setup, where an attacker adopts unauthorized FR models to identify benign users from their publicly shared facial images. This enables the extraction of sensitive information, posing significant privacy risks. Since many well-trained FR models are readily available—either as open-source implementations or through commercial APIs—attackers can easily obtain automated FR tools to achieve this malicious objective. Formally, let $\mathcal{X}$ be the input facial image space and $\mathcal{Y}$ be the set of possible identities. Given a collection of online-scraped face images $\mathcal{D}$, the adversary aims to correctly identify victim users from their facial images according to the following objective:
@@ 2.1 Threat Model
where $x$ stands for a face image, $y$ is the corresponding ground-truth identity, and $\mathrm{FR}:\mathcal{X}\rightarrow\mathcal{Y}$ denotes a FR model.
@@
A typical FR model consists of three key components: a feature extractor, a gallery database, and a query matching step (see [34, 7] for detailed descriptions). Aligned with prior work [26, 30], we assume that FR models employed by the attacker can achieve high accuracy on clean images. Since FR systems can vary in design, the attacker may utilize different feature extractors and query-matching schemes. To comprehensively assess performance under black-box conditions, we evaluate a range of well-trained FR models, from open-source models to commercial APIs. Additionally, we consider face verification as a specialized FR variant, where the gallery database contains only a single user.
2.2 Anti-Facial Recognition
@@
The primary goal of AFR is to evade the identification of unauthorized facial recognition models used by adversaries described in Section 2.1, improving the protection of benign users’ privacy. In this work, we focus on the most popular adversarial-based AFR techniques, which craft imperceptible or naturalistic perturbations (e.g., adversarial makeup) to user’s images to fool FR systems. Specifically, the objective of adversarial-based AFR can be cast into a constrained optimization problem as follows:
@@
$\max\>\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\mathds{1}\big{\{}\mathrm{FR}\big{(}\mathrm{AFR}(x)\big{)}\neq y\big{\}}$ (2)

### v2

# MASQUE: Localized Adversarial Makeup Generation with Text-Guided Diffusion Models

@@ Abstract
As face recognition is increasingly adopted for government and commercial services, its potential misuse has raised serious concerns about privacy and civil rights. To counteract, various anti-facial recognition techniques have been proposed for privacy protection by adversarially perturbing face images, among which generative makeup-based approaches are the most popular. However, these methods, designed primarily to impersonate specific target identities, can only achieve weak dodging success rates while increasing the risk of targeted abuse. In addition, they often introduce global visual artifacts or a lack of adaptability to accommodate diverse makeup prompts, compromising user satisfaction. To address the above limitations, we develop MASQUE, a novel diffusion-based framework that generates localized adversarial makeups guided by user-defined text prompts. Built upon precise null-text inversion, customized cross-attention fusion with masking, and a pairwise adversarial guidance mechanism using images of the same individual, MASQUE achieves robust dodging performance without requiring any external identity. Comprehensive evaluations on open-source facial recognition models and commercial APIs demonstrate that MASQUE significantly improves dodging success rates over all baselines, along with higher perceptual fidelity preservation, stronger adaptability to various makeup prompts, and robustness to image transformations.
@@ 1 Introduction
Facial recognition (FR) systems have been adopted in a wide range of security, biometrics, and commercial applications (Parkhi et al., 2015). However, their unregulated deployment poses privacy risks, allowing for unauthorized surveillance and malicious tracking. To address these concerns, anti-facial recognition (AFR) technologies have emerged to protect user privacy from unauthorized FR systems (Wenger et al., 2023). AFR techniques vary depending on which stage of facial recognition they disrupt and typically work by modifying images before they are shared online. Among them, adversarial methods are particularly effective, subtly altering face images to evade detection while preserving their natural appearance. Traditional adversarial approaches, such as noise-based methods (Joon Oh et al., 2017; Yang et al., 2021; Zhong & Deng, 2022), obscure facial features with norm-bounded global perturbations, while patch-based techniques (Xiao et al., 2021; Komkov & Petiushko, 2021) optimize adversarial patterns in localized image regions. Unfortunately, these methods often introduce noticeable visual artifacts, compromising the usability of the produced face images.
@@
To overcome these limitations, recent work has shifted toward generative approaches, leveraging generative adversarial networks (GANs) or diffusion models for anti-facial recognition. Instead of relying on pixel-wise perturbations, these methods generate unrestricted yet semantically meaningful modifications that blend naturally into facial features. A notable direction is makeup-based AFR (Yin et al., 2021; Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024; Fan et al., 2025), which seamlessly integrates adversarial perturbations into makeup—a plausible approach, as makeup is inherently associated with facial appearance (see Section 2 for detailed discussions of related work).
@@
While makeup-based AFR techniques offer a promising balance between privacy protection and aesthetics, they often struggle to preserve fine-grained facial details or fully adhere to user instructions from diverse prompts. In addition, these methods require images of an external target identity to guide the generation process for adversarial makeup transfer and primarily focus on the impersonation setting for evaluation, increasing the risks of targeted abuse. When considering the more privacy-critical dodging scenarios, their performance in protection success rates often drops significantly.
Table 1: Highlights of the key desirable features of MASQUE compared with prior AFR methods.
Dodging External ID Guidance Localization Prompt-Following Ability
TIP-IM (Yang et al., 2021) - -
AMT-GAN (Hu et al., 2022) image low
C2P (Shamshad et al., 2023) text medium
DiffAM (Sun et al., 2024) image medium
MASQUE (ours) text high
@@
Contributions. Recognizing a few limitations of prior state-of-the-art AFR methods (Section 4.1), we develop MASQUE, a novel diffusion-based image editing framework for localized adversarial makeup generation with customized text guidance (Section 4.2). In particular, MASQUE stands out by achieving the following desiderata simultaneously (see Table 1 for a summary of its key distinctive features):
  • Inoffensive Identity Protection. Built upon pairwise adversarial guidance, MASQUE ensures high dodging success rates without requiring face images of any external identity other than the victim’s. Such a property reduces the risks of targeted misuse and ethical concerns (Section 3.3).
  • Localized Modification. MASQUE employs a facial mask generation and regularization module to constrain adversarial perturbations to designated areas, while preserving fine-grained details of the original face image, thereby vastly enhancing the visual quality of protected images.
  • Strong User Control. By utilizing cross-attention fusion with masking, MASQUE achieves strong prompt-following capability and is adaptable to diverse makeup prompts, providing greater user control and convenience than existing state-of-the-art AFR methods.
Through comprehensive experiments on two popular image benchmarks, a range of FR models and their variants, along with multiple makeup text prompts, MASQUE not only shows significantly higher dodging success rates over baseline methods, but is also capable of preserving visual quality and fine details of original face images (Section 5). In addition, we also demonstrate the robustness of MASQUE’s performance to diverse makeup prompts and various image transformations (Section 6). All of the above suggest the potential of MASQUE as a superior solution to real-world AFR applications.
2 Related Work
Anti-Facial Recognition. Earlier works adopted obfuscation techniques to obscure the facial identity features (Newton et al., 2005) or craft $\ell_{p}$ perturbations to fool FR models (Yang et al., 2021). While effective, these methods often compromise image quality, limiting their real-world applicability. Poisoning-based approaches (Cherepanova et al., 2021; Shan et al., 2020) introduce a new protection scheme by injecting subtle adversarial noise into images to degrade the effectiveness of recognition models. These techniques excel at disrupting model training without visibly altering the image, but rely on strong assumptions about when and how the unauthorized FR model is constructed. Recently, adversarial makeup (Yin et al., 2021; Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024; Fan et al., 2025) has emerged as a promising solution to realize the goal of facial privacy protection. These methods leverage the strong generative capability of generative models to embed adversarial perturbations into natural makeup-based facial modifications, deceiving attackers’ facial recognition models while preserving the aesthetic appeal. For example, Hu et al. (2022) proposed AMT-GAN, which introduces a regularization module and a joint training pipeline for adversarial makeup transfer within the GAN framework (Goodfellow et al., 2014). Advancements in generative models have enhanced the performance of adversarial makeup transfer, such as Clip2Protect (C2P) (Shamshad et al., 2023) and DiffAM (Sun et al., 2024), which adopt a StyleGAN model or a diffusion-based framework, to enable seamless adversarial face modifications with much improved visual quality. In this work, we build upon these developments by leveraging diffusion models to generate visually consistent, localized adversarial makeup for facial privacy protection under dodging scenarios.
More recently, a line of research (Liu et al., 2023; 2024; Wang et al., 2025; Han et al., 2025; Salar et al., 2025) proposed leveraging diffusion models to generate imperceptible AFR-protections without focusing on makeups. For instance, DiffProtect (Liu et al., 2023) uses the diffusion autoencoder model and adversarially guides the semantic code of the original image towards a target face, whereas Salar et al. (2025) improved the prior approach by learning unconditional embeddings as null-text guidance and adversarially modifying the latent code in the latent diffusion model. Our work complements these methods, where we focus on improving the state-of-the-art makeup-based AFR approaches without relying on images from any external identity, while preserving the visual quality of original face images with a strong emphasis on localizing and controlling the added perturbations.
Diffusion-Based Image Editing. Localized editing methods extend diffusion models for targeted modifications. For example, mask-based approaches (Avrahami et al., 2023; Couairon et al., 2022) constrain edits to specific regions using spatial masks, preserving unedited areas but often struggling with structural consistency. Attention-based approaches (Hertz et al., 2022; Tumanyan et al., 2023) guide edits via attention injection, achieving better global structure preservation but suffering from editing leakage. In addition, text-guided diffusion models, capable of synthesizing high-quality images from natural language descriptions, significantly promote the popularization of generative AI. Built on denoising diffusion probabilistic models (Ho et al., 2020), they iteratively refine random noise into coherent images guided by text prompts. Recent advancements in diffusion models have expanded their capabilities to tasks such as localized editing and controllable generation. Mask-based approaches (Avrahami et al., 2023; Couairon et al., 2022) achieve local text-guided modifications by incorporating user-defined constraints, like masks, to confine edits to specific regions. While effective in preserving unedited areas, these methods often struggle to maintain structural consistency within the edited regions, especially in complex scenarios. On the other hand, mask-free attention-based methods (Hertz et al., 2022; Cao et al., 2023; Tumanyan et al., 2023) use attention injection mechanisms to guide edits without requiring explicit masks. These methods excel at preserving the global structure of the image but are prone to editing leakage, where changes unintentionally affect areas beyond the intended region. In this work, we leverage mask-based cross-attention guidance (Mao et al., 2023) to achieve localized adversarial perturbations in the form of makeup. This approach ensures the precise application of modifications to desired regions, guided by user-specified prompts, while addressing the limitations of structural inconsistencies seen in previous methods.
@@ 3 Problem Formulation
3.1 Facial Recognition (FR)
@@
Typically, a FR model consists of three components: a feature extractor, a gallery database, and a query matching scheme (see Wenger et al. (2023) for detailed descriptions). Since FR models can vary in design, the attacker may utilize different feature extractors and query-matching schemes. To comprehensively assess performance under black-box conditions, we evaluate a range of well-trained FR models, from open-source models to commercial APIs. In addition, we consider face verification as a specialized FR variant, where the gallery database contains only a single user’s images.
3.2 Anti-Facial Recognition (AFR)
@@
The goal of AFR is to evade face recognition, thereby improving the protection of benign users’ privacy. We focus on the most popular adversarial-based AFR techniques, which craft imperceptible or naturalistic perturbations (e.g., adversarial makeup) to users’ face images to fool FR models. Specifically, the objective can be cast into a constrained optimization problem detailed as follows:
@@
$\max\>\frac{1}{|\mathcal{D}|}\sum_{(\bm{x},y)\in\mathcal{D}}\mathds{1}\big{\{}\mathrm{FR}\big{(}\mathrm{AFR}(\bm{x})\big{)}\neq y\big{\}},\quad\text{s.t.}\>\>\frac{1}{|\mathcal{D}|}\sum_{(\bm{x},y)\in\mathcal{D}}\Delta\big{(}\mathrm{AFR}(\bm{x}),\bm{x}\big{)}\leq\gamma,$ (1)
where $\mathrm{AFR}:\mathcal{X}\rightarrow\mathcal{X}$ denotes the AFR perturbation function, $\Delta$ is a similarity metric that measures the visual distortion of the AFR-perturbed image $\mathrm{AFR}(\bm{x})$ with reference to the original image $\bm{x}$, and $\gamma>0$ is a threshold parameter reflecting the distortion upper bound that can be tolerated.
As characterized by the optimization objective in Equation 1, a desirable AFR technique is expected to achieve a high dodging success rate (DSR) against facial recognition, which is the primary evaluation metric when comparing the effectiveness of different AFR techniques. For simplicity, we also refer to these users as defenders, who apply AFR techniques to modify their facial images with adversarial perturbations. Since the defender typically lacks precise knowledge of the FR model employed by the attacker, we consider the black-box scenario when evaluating the DSR of each AFR method.
From the defender’s perspective, images perturbed by AFR should appear natural and closely resemble the original, as captured by the optimization constraint in Equation 1. If $\mathrm{AFR}(\bm{x})$ has an unsatisfactory image quality, users may be reluctant to share the protected face image online, even if a high DSR is achieved. Given this expectation of high visual quality, earlier privacy-preserving techniques such as face obfuscation or anonymization (Newton et al., 2005; Sun et al., 2018) are not suitable. In comparison, makeup-based AFR methods (Shamshad et al., 2023; Sun et al., 2024) stand out, as they introduce natural adversarial perturbations with minimal visual distortion. In this work, we aim to improve the performance of generative makeup-based AFR, enhancing both DSR and visual quality.
3.3 Why Dodging instead of Impersonation?
In contrast to the untargeted dodging objective, several state-of-the-art studies (Yang et al., 2021; Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024) have focused on improving AFR performance in targeted impersonation scenarios. Impersonation-oriented methods aim to adversarially modify a user’s facial images to mislead a FR model into recognizing them as a specific target identity. Let $\mathrm{AFR}_{\mathrm{I}}:\mathcal{X}\rightarrow\mathcal{X}$ be the perturbation function with respect to an impersonation-oriented method. In prior work, the impersonation success rate (ISR) is usually computed using the following metric:
$\text{ISR}:=\frac{1}{|\mathcal{D}|}\sum_{(\bm{x},y)\in\mathcal{D}}\mathds{1}\big{\{}\mathrm{FR}\big{(}\mathrm{AFR}_{\mathrm{I}}(\bm{x})\big{)}=y_{t}\},$ (2)
where $y_{t}$ is the target identity with $y_{t}\neq y$. Below, we explain why impersonation is not appropriate.
From the defender’s perspective, impersonation and dodging serve different goals, as reflected in their distinct evaluation metrics (ISR and DSR). As shown by Zhou et al. (2024), high ISR does not guarantee high DSR. If the target identity $y_{t}$ is not included in the attacker’s model (e.g., using facial verification with only a single image of the victim identity $y$), optimizing for higher ISR offers little benefit due to misaligned objectives. As will be shown in Section 1, prior impersonation-oriented AFR methods can only achieve weak protection performance when evaluated in dodging scenarios. In addition, impersonation techniques can lead to a higher likelihood of targeted misuse. Unlike dodging, impersonation-oriented approaches allow users to control the target identity they are mistaken for, making them more susceptible to abuse. This ability can be exploited for malicious purposes, such as falsely associating one’s actions with another individual. For instance, if a user $y$ intentionally applies an AFR tool to impersonate a target $y_{t}$ and engages in illegal activity, FR systems may misidentify $y_{t}$ as the suspect. This deliberate identity manipulation poses significant ethical and legal risks (Wenger et al., 2023). Therefore, effective privacy protection should aim for untargeted dodging success.
Instead of redirecting recognition to a specifically chosen identity, dodging causes misclassification in an uncontrolled manner, typically as an unknown or incorrect identity. This lack of control makes dodging significantly less susceptible to intentional misuse, as users cannot deliberately impersonate or frame a different party. While dodging still prevents correct identification, it does so without assigning a specific false identity, reducing the risk of targeted deception or malicious intent.
4 Proposed Method: MASQUE
4.1 Limitations of Existing AFR Methods
Figure 1: Each row presents the reference image with a text makeup prompt, the original image, followed by AFR-protected images. Methods marked with $\dagger$ require only a text prompt for makeup application and do not rely on a reference image. The yellow text indicates the cosine similarity score. Below the images, we report the averaged impersonation and dodging success rates of each method.
We conduct preliminary experiments to understand the performance of existing AFR methods in terms of privacy protection and visual quality. Figure 1 shows qualitative comparisons of three makeup styles (“red lipstick”, “blue eyebrow”, and “pink eyeshadow”), evaluating several adversarial-based AFR methods, including one of the best current noise-based methods, TIP-IM (Yang et al., 2021), and three state-of-the-art makeup-based approaches (Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024). For each method, we report both the metrics of impersonation and doding success rates across four black-box facial verification models averaged across 100100100 randomly selected images from CelebA-HQ (Karras et al., 2017) and the three makeup styles. To ensure a fair comparison with image-guided AFR methods, we either select the corresponding makeup image from a benchmark facial makeup dataset (Li et al., 2018) or generate one by inpainting the makeup on a non-makeup image (see Section 5.1 and Appendix B for more experimental details of this preliminary study).
Weak Protection under Dodging. As discussed in Section 3.3, an ideal AFR should account for dodging scenarios to ensure privacy without dependence on a target identity. However, despite its ethical advantages, dodging remains an underexplored and underdeveloped area in makeup-based AFR. Existing methods that perform well under impersonation settings often exhibit a significant drop in DSR, as shown in Figure 1. This highlights a fundamental gap: current techniques are not optimized for dodging, leaving privacy protection less reliable when a user seeks to avoid recognition rather than impersonate another identity. In addition, while some AFR methods like Shamshad et al. (2023); Fan et al. (2025) introduce a variant for dodging, they still rely on a target identity. This reliance contradicts the very goal of dodging-based AFR, which should function without any identity substitution. Thus, a key challenge in AFR research is developing a robust dodging-based approach that provides strong protection while avoiding the ethical risks associated with impersonation.
Non-Localized Face Modification. Noise-based approaches (TIP-IM) generate global perturbations, which fail to localize the edits. Makeup-based AFR embeds perturbations within makeup to achieve a more natural disguise, but many GAN-based approaches (AdvMakeup, AMT-GAN, and C2P) introduce global artifacts that alter non-facial regions. This lack of precise localization is particularly problematic when users seek to protect only their facial identity while preserving the background. As shown in Figure 1, this issue is especially pronounced in complex backgrounds, where perturbations should remain confined to facial regions but instead extend beyond. In contrast, DiffAM enables better localization, effectively applying makeup perturbations from a reference image. However, its reliance on reference images limits flexibility, particularly when suitable references are unavailable. This highlights the need for methods that achieve both precise localization and greater flexibility.
Limited Prompt-Following Ability. Makeup-based AFR aim methods aim to modify facial appearance to prevent accurate identification by FR models. However, existing methods offer limited user control, making customization and usability a challenge. Below, we elaborate on this challenge for makeup-based AFR methods based on reference images and text prompts, respectively:
  • Image-reference-based AFR transfers makeup styles from a reference image. While they aim for accurate replication, they struggle with reliability and flexibility. As shown in Figure 1, early GAN-based methods like AMT-GAN fail to transfer makeup consistently, while DiffAM, despite improvements, remains restricted. It only applies makeup to three fixed regions—skin, mouth, and eyes—preventing transfer to other areas like colored eyebrows. In addition, these methods require separate fine-tuning for each reference image, incurring high computational overhead.
4.2 Detailed Design
To address the limitations of existing AFR methods discussed in Section 4.1, we propose MASQUE, a method designed to disrupt FR models with localized adversarial makeup while ensuring no external identity is introduced. Figure 2 illustrates the pipeline of MASQUE with its pseudocode detailed in Algorithm 1 in Appendix A. The makeup generation process is guided by a user-defined text prompt $p^{*}$, refined via cross-attention fusion with a mask. Pairwise adversarial guidance is further introduced to ensure these perturbations mislead FR models without compromising visual fidelity. In particular, our method leverages the state-of-the-art Stable Diffusion (Rombach et al., 2022) adapted to generate high-quality adversarial makeups by iterative denoising a random noise latent ${\bm{z}}_{T}$ conditioned on a text embedding $\bm{c}$. To be more specific, Stable Diffusion is trained to predict the added noise $\bm{\epsilon}$ via:
$\min_{\theta}\ \mathbb{E}_{{\bm{z}}_{0},\bm{\epsilon}\sim\mathcal{N}(\bm{0},\mathrm{I}),t\sim\mathrm{Unif}(1,T)}\bigg{[}\big{\|}\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{z}_{t},t,\bm{c})\big{\|}_{2}^{2}\bigg{]}.$ (3)
Applying Makeup with Cross Attention. Before applying edits, we first obtain a faithful latent representation of the original image $\bm{x}$ using null-text inversion (Mokady et al., 2023), which mitigates reconstruction errors common in direct DDIM inversion (Song et al., 2020). Conditioning on an empty prompt to align the forward and reverse diffusion trajectories, ensuring near-perfect reconstruction of the image’s structure and identity. With this accurate latent representation, we introduce makeup attributes via the text prompt $p^{*}$ by manipulating the cross-attention (CA) layers of the diffusion model (Rombach et al., 2022), which control how spatial features correspond to semantic tokens. At each diffusion step $\tau$, we extract attention maps $A_{\tau}$ (reconstruction) and $A_{\tau}^{*}$ (editing) and blend them: preserving CA values from $A_{\tau}$ for shared tokens to maintain structure, while incorporating values from $A_{\tau}^{*}$ for makeup-specific tokens in $p^{*}$: